OpenStudy (anonymous):

determine the domain and the range of the function: f(x)=x^2+6x

OpenStudy (anonymous):

help

OpenStudy (anonymous):

domain of any polynomial is all real numbers \(\mathbb{R}\)

OpenStudy (anonymous):

range is from the second coordinate of the vertex to \(\infty\) the first coordinate of the vertex is always \(-\frac{b}{2a}\) in your case it is \(-\frac{6}{2\times 1}=-3\) and the second coordinate is what you get when you replace \(x\) by \(-3\)

OpenStudy (anonymous):

would i square the negative 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you would compute \((-3)^2+6\times (-3)\)